# Understanding the Application of Utility Theory in Robotics and Artificial Intelligence: A Survey

### v1

@@ VII Insights and Challenges
Similarly, how to design a general and standard utility mechanism integrating humans and AI agents’ needs and learning to efficiently build reliable and stable trust relationships between agents and humans is essential [128]. Especially a suitable knowledge graph combined with efficient DRL can help agents learn to adopt appropriate behaviors or strategies benefiting the group utilities, adapting to human members’ needs, and guaranteeing their development in various situations [54]. Another interesting direction is behavior manipulation, known as policy elicitation, which refers to problems in human-robot collaboration wherein agents must guide humans toward an optimal policy to fulfill a task through implicit or explicit communication [151, 152]. For example, it is used to achieve effective personalized learning in AI agents’ teaching and coaching [153].
For future MAS research, AI agents learn and adapt to human needs and maintain trust and rapport among agents, which are critical for task efficiency and safety improvement [147]. And we list potential directions as follow:
  • Combining the specific capabilities and needs of agents and humans, calculating sensible strategies to efficiently organize the entire group collaboration and fulfill the corresponding mission.
As discussed above, Agent Needs Hierarchy is the foundation for MAS learning from interaction [55]. It surveys the system’s utility from individual needs. Balancing the needs and rewards between agents and groups for MAS through interaction and adaptation in cooperation optimizes the system’s utility and guarantees sustainable development for each group member, much like human society does. Furthermore, by designing suitable reward mechanisms and developing efficient DRL methods, MAS can effectively represent various group behaviors, skills, and strategies to adapt to uncertain, adversarial, and dynamically changing environments [53].
VIII Conclusion
This paper proposes a utility-orient needs paradigm based on agents’ needs to describe and evaluate inter and outer relationships among agents’ interactions in robotics and AI. We review existing literature illustrating the application of utility theory as a unified concept from the perspective of single-agent, MAS, agent trust, and HRI. Then, we discuss the insight and open problems for future research based on utility theory. We anticipate that this comprehensive overview of the role of the utility-orient needs paradigm based on agents’ needs in robotics and AI systems will support future research that leverages, compares, or constructs interactions for artificial social systems.

### v3

@@ 7.2 Value Systems based Learning
Moreover, how to design a general and standard utility mechanism integrating humans and AI agents’ needs and learning to efficiently build reliable and stable trust relationships between agents and humans is essential [284]. Especially a suitable knowledge graph combined with efficient DRL can help agents learn to adopt appropriate behaviors or strategies benefiting the group utilities, adapting to human members’ needs, and guaranteeing their development in various situations [288]. Another interesting direction is behavior manipulation, known as policy elicitation, which refers to problems in human-robot collaboration wherein agents must guide humans toward an optimal policy to fulfill a task through implicit or explicit communication [248, 54]. For example, it is used to achieve effective personalized learning in AI agents’ teaching and coaching [145].
7.3 Artificial Social Systems
On the other hand, in the development of cognitive modeling in robotics, robots’ cognition will increase, and they will gradually develop self-awareness. They are transforming from traditional roles, like tools and machines, to so-called artificial creatures. Especially with the tremendous growth in AI technology, Robotics, IoT, and high-speed wireless sensor networks (like 5G), it gradually forms an artificial ecosystem termed artificial social systems that involves entire AI agents from software entities to hardware devices. How to integrate artificial social systems into human society and coexist harmoniously is a critical issue for the sustainable development of human beings, such as assistive and healthcare robotics [72], social path planning and navigation [215], search and rescue [193, 283], smart manufacturing [7, 8], and autonomous driving [223]. Especially the future factory is likely to utilize robots for a much broader range of tasks in a much more collaborative manner with humans, which intrinsically requires operation in proximity to humans, raising safety and efficiency issues [249].
Therefore, building a robust, stable, and reliable trust network between humans and AI agents to evaluate agents’ performance and status in a common ground is the pre-condition for efficient and safe interaction in their collaboration. Here, utility is the key concept representing the individual agent’s innate values, needs, and motivations over states of the world and is described as expected rewards. It measures sensitivity to the impact of actions on trust and long-term cooperation and is efficient enough to allow AI agents like robots to make real-time decisions [129]. For example, self-driving cars might be the first widespread case of trustworthy robots designed to earn trust by demonstrating how well they follow social norms. Additionally, research on joint attention [117] between robots, even in human-robot interaction [147], remains a considerable open challenge for modeling social motives such as peer pressure and conformity, especially sharing values or motivations within the group. Moreover, developing an evolutionary value system in cognitive models using utility theory to implement it in cognitive robots, especially for robot swarm interaction [122], will be a more interesting and challenging direction.
Generally speaking, research on cognitive modeling for autonomous robots is still in its infancy, and there is a long way to go before we have a full-fledged autonomous robot with human-like cognition.
8 Conclusion
This survey, from a utility-theoretic perspective, begins with a thorough review of the evolution of cognitive modeling in robotic applications. We discuss cognitive modeling in robotics from BBR and cognitive architectures to the properties of value systems in robots, such as motivations as artificial value systems, and the utility theory based cognitive modeling for generating and updating strategies in robotic interactions. Then, we discussed the applications of biologically inspired motivation systems in single-agent and multi-agent systems. In particular, there has been impressive progress in the design of utility architectures for intrinsic motivation systems implemented in trust evaluation and human-robot interaction. Finally, we analyzed the most important trends of this work to provide a concise representation of the fundamental challenges on which we still need to continue working to deploy these systems in real and lengthy applications.
We believe that cognitive modeling has great potential to accelerate the integration of AI and robotic systems into human society, supporting sustainable human development. This includes alleviating the burden and facilitating the execution of tasks by the most vulnerable sectors of society (such as seniors, children, and the disabled), overcoming the workforce shortage in specific positions related to healthcare, manufacturing, education, etc.